OpenStudy (anonymous):

What is the product written in scientific notation? (2.1 × 10–2)(3.2 × 10–4)

OpenStudy (anonymous):

A. 6.72 × 108 B. 6.72 × 10–6 C. 5.3 × 102 D. 5.3 × 10–6

OpenStudy (anonymous):

\[\large \sf (2.1 \times 3.2) \times (10^{-2} \times 10^{-4})\]

OpenStudy (boldjon):

Multiply the two numbers together. What's 2.1×3.2?

OpenStudy (anonymous):

6.72

OpenStudy (anonymous):

And if they have the same base, you just add the exponents together. \[\large \sf a^{x} \times a^{y} = a^{x+y}\]

OpenStudy (boldjon):

Yep, now add the exponents:

OpenStudy (boldjon):

hold up i'm typing them in

OpenStudy (boldjon):